OpenStudy (anonymous):

The sum of two numbers is x. If one of the numbers is 12, what is the other? 12 - x x - 12 12 + x 12 x

OpenStudy (freckles):

so you are given the sum of two numbers is x. (say the numbers are k and L)-(so L+k=12) one of those numbers is 12. (replace one of the letters there, say replace L with 12) so you are given 12+k=x just solve for other number (which we called k)

OpenStudy (anonymous):

12+x

OpenStudy (freckles):

did you try solving for k ?

OpenStudy (anonymous):

yeah

OpenStudy (freckles):

so how did you solve this equation for k 12+k=x

OpenStudy (anonymous):

12+x just makes sense to me sum of two numbers 12 is one x is the outher

OpenStudy (freckles):

ok since you don't know how to solve the equation above for x I can give you one more way to look at this

OpenStudy (freckles):

you are given them of 12 and another number is x and they are asking what that other number is well 12+which of your choices give you x

OpenStudy (freckles):

choices: 12 - x x - 12 12 + x 12 x --- well is 12+(12-x) = x? or is 12+(x-12)=x? or is 12+(12+x)=x? or is 12+(12x)=x?

OpenStudy (freckles):

which of these equations are true

OpenStudy (freckles):

you are given the sum of 12 and another number is x *

OpenStudy (anonymous):

12x

OpenStudy (freckles):

12x+12 how does this simplify to x?

OpenStudy (anonymous):

idk im not good at algerbra

OpenStudy (freckles):

you don't have to be you have to be good at arithmetic for this question do you know what 12-12=?

OpenStudy (freckles):

12-12... if you have 12 apples and someone takes all 12 apples away then how many apples are you left with?

OpenStudy (anonymous):

0

OpenStudy (freckles):

so what does this simplify to then: 12+(x-12)

OpenStudy (freckles):

you can rearrange it... 12-12+x

OpenStudy (anonymous):

x-12

OpenStudy (freckles):

12-12+x = 0+x = x

OpenStudy (freckles):

and above solving this equation for k : 12+k=x you would subtract 12 on both sides -12 -12 --------------- k=x-12 just going back to how to do the previous way